m^4 - 10m^2 + 9 = 0
Answer by stanbon(48544)

m^4 - 10m^2 + 9 = 0
----------
(m^2-9)(m^2+9) = 0
(m-3)(m+3)(m+3i)(m-3i)=0
m=3 or m=-3 or m=-3i or m=3i
=============
Cheers,
Stan H.